OpenStudy (anonymous):

if y= 5x/x^3 -4, then dy/dx =?

OpenStudy (anonymous):

y=5x/x^3 -4 y=5/x^2-4 y=5*[x^(-2)}-4 dy/dx=5*(-2) [ x^(-2-1)]-0 dy/dx=-10*x^(-3)

OpenStudy (anonymous):

why in the second step the exponent became 2...did you subtract

OpenStudy (anonymous):

dividing x by x^3

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcum ! :-)

OpenStudy (experimentx):

is your expression on right hand side 5x/(x^3 -4) ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well i thought 4 is not included , and acc I have diffrenctd , whereas had it been with x^3 , ans would have been diff

OpenStudy (anonymous):

now if its included , quotient rule has to be applied

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

in the denominator it is....x^(3) - 4

OpenStudy (anonymous):

oops dn my ans is wrong ! wait I'll explain you this one

OpenStudy (anonymous):

\[ f(x) = \frac {5x}{x^3 -4} \implies f'(x) = \frac{(x^3 -4)\times 5 -5x \times(3x)}{(x^3 -4)^2} \]

OpenStudy (anonymous):

quotient rule : {(dy/dx numerator)*denominator-[(dy/dx denominator)*numerator]} / denominator^2 try solving with this , if you get stuck , letme know ! btw FFM has done a great job already , so you wont find any pro ! :-)

OpenStudy (anonymous):

fool for math, why isnt it 3x^2 in the numerator

OpenStudy (anonymous):

Lol, because It was a typo.

OpenStudy (anonymous):

\[ f(x) = \frac {5x}{x^3 -4} \implies f'(x) \] \[= \frac{(x^3 -4)\times 5 -5x \times(3x^2)}{(x^3 -4)^2} = -\frac{10 \left(2+x^3\right)}{\left(-4+x^3\right)^2} \]

OpenStudy (anonymous):

Do check my algebra, don't ever forget that I am a Fool ;) :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

looks right...-10x^3 - 20/ (x^3 - 4)^2

OpenStudy (anonymous):

yea thats right ! :-)

OpenStudy (anonymous):

Wow! Congratz too me :D Yay!!!

OpenStudy (anonymous):

just strict to the formula , and you ll never be wrong !

OpenStudy (anonymous):

She mean stick.

OpenStudy (anonymous):

oops ! typing error , stick ! :p

OpenStudy (anonymous):

lol, it happens to non-fools too :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

LOL ! :D